# Conformal Reconfigurable Intelligent Surfaces: A Cylindrical Geometry Perspective

### v1

@@ III Conclusion
Future work will explore several directions, including higher-bit reconfigurable designs and the incorporation of space-time coding strategies, originally developed for planar RISs [Zhang:2018st], into conformal architectures for enhanced spatial and spectral control. On the experimental side, we are progressing toward prototype fabrication and validation, with laboratory measurements to be followed by field tests on UAV platforms within a dedicated flight arena under construction at the Italian Aerospace Research Centre.
@@ IV-A Numerical Simulations
Full-wave simulations are performed using the finite-element-based commercial software Ansys HFSS [HFSS]. Two types of simulations are carried out: meta-atom characterization and full-structure analysis.
@@ IV-A1 Meta-atom characterization
The meta-atom geometry used for characterization is shown in the inset of Figure 7. The square metallic patch has side length $L=26$ mm and is printed on a grounded substrate of thickness $h=1.57$ mm, with a lattice period of $p=38$ mm. The substrate is modeled using its nominal material parameters, namely a relative permittivity $\varepsilon_{r}=2.2$ and a loss tangent $\tan\delta=0.001$. The patch is connected to the ground plane through a metallic via incorporating a PIN diode, which enables binary reconfiguration. The diode is modeled as a lumped-element series circuit, consisting of an RL pair (1.5 $\Omega$, 0.7 nH) in the ON state and an LC pair (0.7 nH, 0.15 pF) in the OFF state [Yang:2016ap]. All metals are treated as perfect electric conductors. To emulate an infinitely periodic array, the unit cell is truncated along the transverse directions using phase-shift (Floquet) boundary conditions.
@@ IV-B Optimization
The ES method minimizes the objective function in Equation (14) by exhaustively searching all possible combinations of PIN diode states across the RIS. The GA-based optimization is carried out using an in-house code implemented in MATLAB [MATLAB], with the goal of minimizing the maximum field magnitude outside the desired angular window. The GA is configured with a population size of 1000, 200 generations, a crossover probability of 0.9, and a mutation probability of 0.05. The MPDR method, which provides a closed-form low-complexity solution, is described in detail in the Appendix.
@@ Appendix A Details on GO Synthesis
Following the approach used for planar anomalous reflectors [Sun:2012he], the local reflection coefficient can be approximated as
@@
Alternatively, the reflection coefficient can be expressed in terms of the surface impedance as [Sun:2012he]
@@
where the wave impedance of the incident field is [Harrington:2001th]
@@ Appendix B Details on MPDR Synthesis
To reduce the computational complexity of the synthesis process, we propose a formulation based on the MPDR beamforming strategy [Cap1969]. The problem is cast as a discrete optimization task:
@@
By relaxing the constraint in Equation (19c) and allowing $\boldsymbol{\gamma}\in\mathbb{C}^{N}$, the problem in (19) becomes a linearly constrained quadratic programming problem. This type of problem admits a closed-form solution [Cap1969], given by

### v2

@@ III Conclusion
Future work will explore several directions, including higher-bit reconfigurable designs and the incorporation of space-time coding strategies, originally developed for planar RISs [49], into conformal architectures for enhanced spatial and spectral control. On the experimental side, we are progressing toward prototype fabrication and validation, with laboratory measurements to be followed by field tests on UAV platforms within a dedicated flight arena under construction at the Italian Aerospace Research Centre.
@@ IV-A Numerical Simulations
Full-wave simulations are performed using the finite-element-based commercial software Ansys HFSS [3]. Two types of simulations are carried out: meta-atom characterization and full-structure analysis.
@@ IV-A1 Meta-atom characterization
The meta-atom geometry used for characterization is shown in the inset of Figure 7. The square metallic patch has side length $L=26$ mm and is printed on a grounded substrate of thickness $h=1.57$ mm, with a lattice period of $p=38$ mm. The substrate is modeled using its nominal material parameters, namely a relative permittivity $\varepsilon_{r}=2.2$ and a loss tangent $\tan\delta=0.001$. The patch is connected to the ground plane through a metallic via incorporating a PIN diode, which enables binary reconfiguration. The diode is modeled as a lumped-element series circuit, consisting of an RL pair (1.5 $\Omega$, 0.7 nH) in the ON state and an LC pair (0.7 nH, 0.15 pF) in the OFF state [47]. All metals are treated as perfect electric conductors. To emulate an infinitely periodic array, the unit cell is truncated along the transverse directions using phase-shift (Floquet) boundary conditions.
@@ IV-B Optimization
The ES method minimizes the objective function in Equation (14) by exhaustively searching all possible combinations of PIN diode states across the RIS. The GA-based optimization is carried out using an in-house code implemented in MATLAB [34], with the goal of minimizing the maximum field magnitude outside the desired angular window. The GA is configured with a population size of 1000, 200 generations, a crossover probability of 0.9, and a mutation probability of 0.05. The MPDR method, which provides a closed-form low-complexity solution, is described in detail in the Appendix.
@@ Appendix A Details on GO Synthesis
Following the approach used for planar anomalous reflectors [44], the local reflection coefficient can be approximated as
@@
Alternatively, the reflection coefficient can be expressed in terms of the surface impedance as [44]
@@
where the wave impedance of the incident field is [18]
@@ Appendix B Details on MPDR Synthesis
To reduce the computational complexity of the synthesis process, we propose a formulation based on the MPDR beamforming strategy [7]. The problem is cast as a discrete optimization task:
@@
By relaxing the constraint in Equation (19c) and allowing $\boldsymbol{\gamma}\in\mathbb{C}^{N}$, the problem in (19) becomes a linearly constrained quadratic programming problem. This type of problem admits a closed-form solution [7], given by